How do you factor the trinomial: 3(x+2)^2+12(x+2)-15
Thank you!

Answer by edjones(8007)

3(x+2)^2+12(x+2)-15
=3(x^2+4x+4)+12x+24-15
=3x^2+12x+12+12x+9
=3x^2+24x+21
=3(x^2+8x+7)
=3(x+1)(x+7)
.
Ed